OpenStudy (anonymous):

Jade hiked 1 1/4 mile on the trail. Gretta hiked 1 1/2 times as far on the trail as Jade. How far did Gretta hike on the trail?

OpenStudy (theraggedydoctor):

Multiply \(1\frac{ 1 }{ 4 } \times 1 \frac{ 1 }{ 2 }\)

OpenStudy (hockeychick23):

so for this we do 5/4*3/2 which is 15/8 which simplified is 1 7/8 miles which could also be written as 1.875 miles